OpenStudy (anonymous):

Will Medal And Fan Best Answer!! I need explanation AND answer! Suppose that 12 inches of wire costs 48 cents. At the same rate, how many inches of wire can be bought for 28 cents?

OpenStudy (anonymous):

\[\large \sf 12 \times cost~per~inch~=~48\] So what is the cost per inch?

OpenStudy (anonymous):

48 / 12 = 4 Cost per inch is 4 cents

OpenStudy (anonymous):

I'm not asking you

OpenStudy (anonymous):

28 / 4 = 7 7 inches can be bought

OpenStudy (anonymous):

7 inches lol